OpenStudy (anonymous):

Let t be the time now. Write an expression for the time 2 hours ago.

OpenStudy (anonymous):

\[t-2\]

jimthompson5910 (jim_thompson5910):

Expression is t-2 hours

jimthompson5910 (jim_thompson5910):

Ex: say t = 10 hours, then t-2 = 10 - 2 = 8 hours

OpenStudy (anonymous):

suppose it is 1?

OpenStudy (anonymous):

\[t-2\equiv (\text{ mod }12)\]

jimthompson5910 (jim_thompson5910):

hmm oh I meant 10 as in 10 o clock lol, oh well it still applies; however, you have to keep in mind that the time will wrap around

OpenStudy (anonymous):

2hours ago =120m there fore ct (current time) -120pt (past time)